OpenStudy (purplemexican):

What is the probability of getting tails each time if a coin is tossed 4 times?

OpenStudy (purplemexican):

1/256 1/16 1/8 2

OpenStudy (anonymous):

well i mean idk what he is sayin g but uhh

OpenStudy (purplemexican):

1/2*1/2*1/2*1/2 is what im trying to get

Mehek (mehek14):

\(\dfrac{1}{2}*\dfrac{1}{2}*\dfrac{1}{2}*\dfrac{1}{2}\)

OpenStudy (purplemexican):

i get 0.062 or 31/500 but that's not a choice

Mehek (mehek14):



Mehek (mehek14):

obviously 1*1*1*1 = 1 so the numerator is 1

OpenStudy (purplemexican):

thank you

Mehek (mehek14):

to find the denominator \(2^4\)

Mehek (mehek14):

yw ^_^